# On forbidden configurations in

### v1

@@ 5 Proof of Theorem 4
Thus, we assume that all 4-tuples of black vertices from $H_{3}$ are colored with $\pi$ for some permutation $\pi\in S_{4}$. We show that all these cases are impossible. Let $b^{\prime}_{1},\dots,b^{\prime}_{2k}$ be the vertices of $H_{3}$ such that the label of $b^{\prime}_{i}$ in $H$ is smaller than the label of $b^{\prime}_{j}$ in $H$ if and only if $i<j$. First, we observe that, since $2k\geq 6$, either $\pi(1)$ and $\pi(2)$ are consecutive in $<$ or $\pi(3)$ and $\pi(4)$ are consecutive in $<$. If, for example, $\pi(1)<\pi(3)<\pi(2)<\pi(4)$, then by considering 4-tuples $(b^{\prime}_{1},b^{\prime}_{3},b^{\prime}_{5},b^{\prime}_{6})$ and $(b^{\prime}_{1},b^{\prime}_{4},b^{\prime}_{5},b^{\prime}_{6})$, we see that the line containing the edge $b^{\prime}_{1}w(b^{\prime}_{1},b^{\prime}_{5})$ also contains white points $w(b^{\prime}_{3},b^{\prime}_{6})$ and $w(b^{\prime}_{4},b^{\prime}_{6})$. On the other hand, by considering the 4-tuples $(b^{\prime}_{2},b^{\prime}_{3},b^{\prime}_{5},b^{\prime}_{6})$ and $(b^{\prime}_{2},b^{\prime}_{4},b^{\prime}_{5},b^{\prime}_{6})$, the line containing the edge $b^{\prime}_{2}w(b^{\prime}_{2},b^{\prime}_{5})$ also contains white points $w(b^{\prime}_{3},b^{\prime}_{6})$ and $w(b^{\prime}_{4},b^{\prime}_{6})$. This is impossible as we then have a line containing an edge of $H_{3}$ and two black points of $H_{3}$. The other cases are analogous. In fact, $\pi(3)$ and $\pi(4)$ have to be consecutive in $<$. If, for example, $\pi(3)<\pi(1)<\pi(2)<\pi(4)$, then the 4-tuples $(b^{\prime}_{1},b^{\prime}_{2},b^{\prime}_{4},b^{\prime}_{5})$, $(b^{\prime}_{1},b^{\prime}_{2},b^{\prime}_{4},b^{\prime}_{6})$, $(b^{\prime}_{1},b^{\prime}_{3},b^{\prime}_{4},b^{\prime}_{5})$, and $(b^{\prime}_{1},b^{\prime}_{3},b^{\prime}_{4},b^{\prime}_{6})$ similarly as before imply that there is a line containing $b^{\prime}_{2}w(b^{\prime}_{2}b^{\prime}_{4})$ and $b^{\prime}_{3}w(b^{\prime}_{3}b^{\prime}_{4})$, which is forbidden in $H_{3}$. The remaining cases are again analogous. Similarly, if $\pi(1)$ and $\pi(2)$ are consecutive, but $\pi(3)$ and $\pi(4)$ are not, then we have, say, $\pi(3)<\pi(1)<\pi(2)<\pi(4)$. If we consider the 4-tuples $(b^{\prime}_{1},b^{\prime}_{3},b^{\prime}_{5},b^{\prime}_{6})$, $(b^{\prime}_{2},b^{\prime}_{3},b^{\prime}_{5},b^{\prime}_{6})$, $(b^{\prime}_{1},b^{\prime}_{4},b^{\prime}_{5},b^{\prime}_{6})$, and $(b^{\prime}_{2},b^{\prime}_{4},b^{\prime}_{5},b^{\prime}_{6})$, then we obtain a line that contains $b^{\prime}_{3}w(b^{\prime}_{3}b^{\prime}_{5})$ and $b^{\prime}_{4}w(b^{\prime}_{4}b^{\prime}_{5})$, which is impossible. Finally, we show that the last case when both $\pi(1)$ and $\pi(2)$ are consecutive in $<$ as well as $\pi(3)$ and $\pi(4)$ is impossible as well. For example, if $\pi(1)<\pi(2)<\pi(3)<\pi(4)$, then it suffices to consider the 4-tuples $(b^{\prime}_{1},b^{\prime}_{2},b^{\prime}_{4},b^{\prime}_{5})$, $(b^{\prime}_{1},b^{\prime}_{2},b^{\prime}_{4},b^{\prime}_{6})$, $(b^{\prime}_{2},b^{\prime}_{3},b^{\prime}_{4},b^{\prime}_{5})$, and $(b^{\prime}_{2},b^{\prime}_{3},b^{\prime}_{4},b^{\prime}_{6})$ and find a line containing $b^{\prime}_{1}w(b^{\prime}_{1}b^{\prime}_{2})$ and $b^{\prime}_{2}w(b^{\prime}_{2}b^{\prime}_{3})$, which is forbidden in $H_{3}$.
@@
Since $H$ is a 1-subdivision of $K_{R^{\prime}}$, it is a bipartite graph with maximum degree 2 on one side. Thus, by Theorem 18, we have $ex(m,H)\in O(m^{3/2})$. Recall that $G$ contains at least $\Omega(m\log^{2}{m})$ edges. Therefore, using Theorem 17 for $\alpha=1/2$, we get that
$\textrm{cr}(G)\geq c^{\prime}\frac{(I-n)^{2+2}}{m^{1+2}}\geq c^{\prime}\frac{(I/4)^{4}}{m^{3}}$
for some constant $c^{\prime}=c^{\prime}(k)>0$. On the other hand, since any two lines intersect at most once, we also have $\textrm{cr}(G)\leq\binom{n}{2}$. Combining these upper and lower bounds on $\textrm{cr}(G)$ implies the result. ∎
@@ 6 Proof of Theorem 6
By Markov’s inequality, we have $X\leq\frac{qN}{64}$ with probability at least $\frac{1}{2}$. Then, there exists $B^{\prime}$ such that $\mathcal{A}$ holds and the number of copies of subdivided $k$-cliques in $B^{\prime}$ is at most $\frac{qN}{64}$. Delete a vertex from each copy, and let $B^{\prime\prime}$ be the resulting graph with parts $P^{\prime\prime}$ and $\mathcal{L}^{\prime\prime}$. This way we delete at most $\frac{qN}{64}$ vertices and at most $\frac{qN}{64}\cdot 2qN^{1/3}$ edges. Then, $B^{\prime\prime}$ has at least $\frac{q^{2}N^{4/3}}{16}-\frac{q^{2}N^{4/3}}{32}\geq\frac{q^{2}N^{4/3}}{32}$ edges and no copy of the incidence graph of a subdivided $k$-clique.
If we choose $n$ such hat $\frac{qN}{8}\leq n\leq\frac{qN}{4}$, then $n=N^{1-\frac{k+6}{9k+6}+o(1)}$. By sampling random $n$ element subsets of $P^{\prime\prime}$ and $\mathcal{L}^{\prime\prime}$, we get that there exists an induced subgraph $B$ of $B^{\prime\prime}$ with parts $P$ and $\mathcal{L}$, each of size $n$, such that $B$ has at least

### v2

@@ 5 Proof of Theorem 4
Thus, we assume that all 4-tuples of black vertices from $H_{3}$ are colored with $\pi$ for some permutation $\pi\in S_{4}$. We show that all these cases are impossible. Let $b^{\prime}_{1},\dots,b^{\prime}_{2k}$ be the vertices of $H_{3}$ such that the label of $b^{\prime}_{i}$ in $H$ is smaller than the label of $b^{\prime}_{j}$ in $H$ if and only if $i<j$. First, we observe that, since $2k\geq 6$, either $\pi(1)$ and $\pi(2)$ are consecutive in << or $\pi(3)$ and $\pi(4)$ are consecutive in <<. If, for example, $\pi(1)<\pi(3)<\pi(2)<\pi(4)$, then by considering 4-tuples $(b^{\prime}_{1},b^{\prime}_{3},b^{\prime}_{5},b^{\prime}_{6})$ and $(b^{\prime}_{1},b^{\prime}_{4},b^{\prime}_{5},b^{\prime}_{6})$, we see that the line containing the edge $b^{\prime}_{1}w(b^{\prime}_{1},b^{\prime}_{5})$ also contains white points $w(b^{\prime}_{3},b^{\prime}_{6})$ and $w(b^{\prime}_{4},b^{\prime}_{6})$. On the other hand, by considering the 4-tuples $(b^{\prime}_{2},b^{\prime}_{3},b^{\prime}_{5},b^{\prime}_{6})$ and $(b^{\prime}_{2},b^{\prime}_{4},b^{\prime}_{5},b^{\prime}_{6})$, the line containing the edge $b^{\prime}_{2}w(b^{\prime}_{2},b^{\prime}_{5})$ also contains white points $w(b^{\prime}_{3},b^{\prime}_{6})$ and $w(b^{\prime}_{4},b^{\prime}_{6})$. This is impossible as we then have a line containing an edge of $H_{3}$ and two black points of $H_{3}$. The other cases are analogous. In fact, $\pi(3)$ and $\pi(4)$ have to be consecutive in <<. If, for example, $\pi(3)<\pi(1)<\pi(2)<\pi(4)$, then the 4-tuples $(b^{\prime}_{1},b^{\prime}_{2},b^{\prime}_{4},b^{\prime}_{5})$, $(b^{\prime}_{1},b^{\prime}_{2},b^{\prime}_{4},b^{\prime}_{6})$, $(b^{\prime}_{1},b^{\prime}_{3},b^{\prime}_{4},b^{\prime}_{5})$, and $(b^{\prime}_{1},b^{\prime}_{3},b^{\prime}_{4},b^{\prime}_{6})$ similarly as before imply that there is a line containing $b^{\prime}_{2}w(b^{\prime}_{2}b^{\prime}_{4})$ and $b^{\prime}_{3}w(b^{\prime}_{3}b^{\prime}_{4})$, which is forbidden in $H_{3}$. The remaining cases are again analogous. Similarly, if $\pi(1)$ and $\pi(2)$ are consecutive, but $\pi(3)$ and $\pi(4)$ are not, then we have, say, $\pi(3)<\pi(1)<\pi(2)<\pi(4)$. If we consider the 4-tuples $(b^{\prime}_{1},b^{\prime}_{3},b^{\prime}_{5},b^{\prime}_{6})$, $(b^{\prime}_{2},b^{\prime}_{3},b^{\prime}_{5},b^{\prime}_{6})$, $(b^{\prime}_{1},b^{\prime}_{4},b^{\prime}_{5},b^{\prime}_{6})$, and $(b^{\prime}_{2},b^{\prime}_{4},b^{\prime}_{5},b^{\prime}_{6})$, then we obtain a line that contains $b^{\prime}_{3}w(b^{\prime}_{3}b^{\prime}_{5})$ and $b^{\prime}_{4}w(b^{\prime}_{4}b^{\prime}_{5})$, which is impossible. Finally, we show that the last case when both $\pi(1)$ and $\pi(2)$ are consecutive in << as well as $\pi(3)$ and $\pi(4)$ is impossible as well. For example, if $\pi(1)<\pi(2)<\pi(3)<\pi(4)$, then it suffices to consider the 4-tuples $(b^{\prime}_{1},b^{\prime}_{2},b^{\prime}_{4},b^{\prime}_{5})$, $(b^{\prime}_{1},b^{\prime}_{2},b^{\prime}_{4},b^{\prime}_{6})$, $(b^{\prime}_{2},b^{\prime}_{3},b^{\prime}_{4},b^{\prime}_{5})$, and $(b^{\prime}_{2},b^{\prime}_{3},b^{\prime}_{4},b^{\prime}_{6})$ and find a line containing $b^{\prime}_{1}w(b^{\prime}_{1}b^{\prime}_{2})$ and $b^{\prime}_{2}w(b^{\prime}_{2}b^{\prime}_{3})$, which is forbidden in $H_{3}$.
@@
Since $H$ is a 1-subdivision of $K_{R^{\prime}}$, Theorem 18 implies that $ex(m,H)\in O(m^{3/2-1/(4R^{\prime}-6)})$. Recall that $G$ contains at least $\Omega(m\log^{2}{m})$ edges. Therefore, using Theorem 17 for $\alpha=1/2-1/(4R^{\prime}-6)$, we get that
$\textrm{cr}(G)\geq c^{\prime}\frac{(I-n)^{2+\frac{1}{1/2-1/(4R^{\prime}-6)}}}{m^{1+\frac{1}{1/2-1/(4R^{\prime}-6)}}}=c^{\prime}\frac{(I-n)^{\frac{4R^{\prime}-7}{R^{\prime}-2}}}{m^{\frac{3R^{\prime}-5}{R^{\prime}-2}}}\geq c^{\prime}\frac{(I/4)^{\frac{4R^{\prime}-7}{R^{\prime}-2}}}{m^{\frac{3R^{\prime}-5}{R^{\prime}-2}}}$
for some constant $c^{\prime}=c^{\prime}(k)>0$. On the other hand, since any two lines intersect at most once, we also have $\textrm{cr}(G)\leq\binom{n}{2}$. Combining these upper and lower bounds on $\textrm{cr}(G)$ gives
$I\in O\left(n^{\frac{2R^{\prime}-4}{4R^{\prime}-7}}m^{\frac{3R^{\prime}-5}{4R^{\prime}-7}}\right)=O\left(n^{1/2-\frac{1}{8R^{\prime}-14}}m^{3/4-\frac{1}{16R^{\prime}-28}}\right),$
which implies the result. ∎
@@ 6 Proof of Theorem 6
By Markov’s inequality, we have $X\leq\frac{qN}{64}$ with probability at least $\frac{1}{2}$. Then, there exists $B^{\prime}$ such that $\mathcal{A}$ holds and the number of copies of subdivided $k$-cliques in $B^{\prime}$ is at most $\frac{qN}{64}$. Delete a vertex from each copy and let $B^{\prime\prime}$ be the resulting graph with parts $P^{\prime\prime}$ and $\mathcal{L}^{\prime\prime}$. In this way, we delete at most $\frac{qN}{64}$ vertices and at most $\frac{qN}{64}\cdot 2qN^{1/3}$ edges. Then, $B^{\prime\prime}$ has at least $\frac{q^{2}N^{4/3}}{16}-\frac{q^{2}N^{4/3}}{32}\geq\frac{q^{2}N^{4/3}}{32}$ edges and no copy of the incidence graph of a subdivided $k$-clique.
If we choose $n$ such hat $\frac{qN}{8}\leq n\leq\frac{qN}{4}$, then $n=N^{1-\frac{k+6}{9k+6}+o(1)}$. By sampling random $n$-element subsets of $P^{\prime\prime}$ and $\mathcal{L}^{\prime\prime}$, we get that there exists an induced subgraph $B$ of $B^{\prime\prime}$ with parts $P$ and $\mathcal{L}$, each of size $n$, such that $B$ has at least